OpenStudy (anonymous):

It would cost 79.23 to buy 3 tickets in Japan and 2 tickets in switzerland. Three tickets in Switzerland and 2 tickets in Japan would cost 75.57. What is the cost of a ticket in each country?

OpenStudy (goformit100):

do you how to form linear equation i two varibles

OpenStudy (anonymous):

that would be 3x+2y=79.23 and 3y+2x=75.57 is this right so far?

OpenStudy (anonymous):

Yes you car right so far. You can use substitution or elimination solve this.

OpenStudy (anonymous):

ok, let me see what i can come up with.

OpenStudy (anonymous):

17.31 and 13.65 is what i get.

OpenStudy (anonymous):

Yes, that's what I got as well...13.65 for y and 17.31 for x

OpenStudy (anonymous):

79.23-75.57=3.66 3.66*3=10.98 75.57+10.98=86.55 86.55/5=17.31 17.31-3.66=13.65

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No prob. I'm not exactly sure how your steps worked, but if you got the same answer I guess your method works as well.